OpenStudy (anonymous):

I was looking at the examples for solving Critical Points and I understand them but the question that I'm working on is ------> f(x)= x^4-2x, find the critical points and determine whether they are Max/Min/Neither. Assistance please, I tried to factor them down and but I'm not sure if my answer is correct or if I even did it properly!? I would appreciate any help, thanks :) p.s I'm a newbie to this 'Open Study' :)

OpenStudy (anonymous):

ok first of all we have a polynomial of degree 4 with positive leading coefficient, so before we start we know it must look something like this 

OpenStudy (anonymous):

taking the derivative we get \[f'(x)=4x^3-2\] and we want the critical points. this one only seems to have one because if you set \[4x^3-2=0\] you get \[4x^3=2\] \[x^3=\frac{1}{2}\] \[x=\sqrt[3]{\frac{1}{2}}\] as the only critical point

OpenStudy (anonymous):

Should I differentiate the function? f(x)= x^4-2x

OpenStudy (anonymous):

yes. that is the first step you get \[f'(x)=4x^3-2\]

OpenStudy (anonymous):

and as you can see this only has one zero, so there is only one critical point. it has to be a minimum for sure because we know what the choices are for a 4th degree polynomial. i will draw the possibilites

OpenStudy (anonymous):

okay thanks so much I forgot to differentiate! Brilliant

OpenStudy (anonymous):

something like these